# GraphXForm: Graph transformer for computer-aided molecular design with application to extraction

### v1

@@ Abstract
Generative deep learning has become pivotal in molecular design for drug discovery and materials science. A widely used paradigm is to pretrain neural networks on string representations of molecules and fine-tune them using reinforcement learning on specific objectives. However, string-based models face challenges in ensuring chemical validity and enforcing structural constraints like the presence of specific substructures. We propose to instead combine graph-based molecular representations, which can naturally ensure chemical validity, with transformer architectures, which are highly expressive and capable of modeling long-range dependencies between atoms. Our approach iteratively modifies a molecular graph by adding atoms and bonds, which ensures chemical validity and facilitates the incorporation of structural constraints. We present GraphXForm, a decoder-only graph transformer architecture, which is pretrained on existing compounds and then fine-tuned using a new training algorithm that combines elements of the deep cross-entropy method with self-improvement learning from language modeling, allowing stable fine-tuning of deep transformers with many layers. We evaluate GraphXForm on two solvent design tasks for liquid-liquid extraction, showing that it outperforms four state-of-the-art molecular design techniques, while it can flexibly enforce structural constraints or initiate the design from existing molecular structures.
Machine Learning, ICML
@@ 1 Introduction
To address the challenge of capturing long-range dependencies in sequential data that RNNs face, transformers (Vaswani, 2017) have been successfully applied due to their ability to model long-range dependencies more efficiently and as they are widely used in language modeling today (Xu et al., 2024). However, transformers are resource-intensive, and needing to fine-tune them with RL can limit the size that is used in practice (Xu et al., 2024; Henderson et al., 2018; Schulman et al., 2017). In general, chemical language models may propose string representations of molecules with invalid chemical structures – for example, when SMILES syntax or valence constraints are violated – which has led to numerous works aimed at circumventing this problem (O’Boyle & Dalke, 2018; Krenn et al., 2020; Cheng et al., 2023; Dai et al., 2018). Despite recent evidence that invalid SMILES can actually be beneficial from a language modeling perspective (Skinnider, 2024), they can harm the RL component of the pipeline as they can increase sample complexity and necessitate reward shaping to account for them. Additionally, the sequential nature of string synthesis makes it difficult to enforce structural constraints, such as ensuring the presence of a minimum number of specific atom types, restrictions on bonds, or even enforcing the presence of certain substructures within the molecules.
@@
An alternative approach is to represent molecules directly as graphs, where atoms are nodes and bonds are edges, and to develop models that modify a molecular graph or design it directly (Wang et al., 2023; Li et al., 2018; Jin et al., 2018; Zhou et al., 2019; Jensen, 2019; De Cao & Kipf, 2018; Zang & Wang, 2020; Zhang et al., 2023b; Mahmood et al., 2021; Maziarka et al., 2019; Rittig et al., 2023d). Working at the graph level, allows explicit encoding of atomic interactions and bonding rules, ensuring chemical validity, and makes it straightforward to start from existing structures and modify them. For example, graph-based methods like GraphGA (Jensen, 2019) employ genetic algorithms to modify molecular graphs directly, even outperforming several neural-based string-synthesis methods without relying on neural networks. As a deep learning example, Zhou et al. (2019) use deep Q-Learning (Mnih, 2013) with a simple feedforward network to optimize graphs from scratch by adding or removing atoms and bonds.
@@
We propose to combine and extend the strengths of both paradigms: employing transformers for their ability to capture long-range dependencies in sequence data, and leveraging pretraining on existing molecules, all while working directly on the molecular graph. More specifically, we use graph transformers (Ying et al., 2021; Maziarka et al., 2019) and formulate molecule design as a sequential task, where a molecular graph – starting from an arbitrary structure – is iteratively extended by placing atoms, bonding them to existing atoms, and modifying the graph by increasing bond orders.
@@
To this end, we introduce GraphXForm, a decoder-only graph transformer architecture that guides these incremental decisions, predicting the next action based on the current molecular graph. Pretrained on existing compounds, GraphXForm employs a novel fine-tuning approach that combines elements of the deep cross-entropy method (Wagner, 2021) and self-improvement learning (Huang et al., 2023; Pirnay & Grimm, 2024a). Unlike commonly used deep reinforcement learning methods like REINFORCE (Williams, 1992), this approach allows for stable training of deep transformers with many layers.
@@
We test GraphXForm for the design of solvents, which are highly relevant for chemical processes. While algorithmic advances in molecular design have primarily focused on de novo drug development, other areas such as materials science and chemical engineering have only recently gained attention. Recent applications include catalyst (Schilter et al., 2023), fuel (Rittig et al., 2023d; Sarathy & Eraqi, 2024), polymer (Jiang et al., 2024; Yue et al., 2024), surfactant (Nnadili et al., 2024), chemical reaction substrate (Nigam et al., 2023), and solvent (König-Mattern et al., 2024) design. We focus specifically on solvents due to their vital role in industrial chemical processes such as reactions, separations, and extractions. While König-Mattern (König-Mattern et al., 2024) recently applied a graph-based genetic algorithm for solvent design, the use of generative ML-based approaches remains underexplored. Therefore, our goal is to compare newer design methods and expand their capabilities.
@@
We evaluate GraphXForm on two liquid-liquid extraction tasks. The objective function in these tasks is defined by a separation factor based on activity coefficients at infinite dilution. We compare GraphXForm on these two downstream tasks to four state-of-the-art molecular design approaches (Graph GA (Jensen, 2019), REINVENT-Transformer (Xu et al., 2024), Junction Tree VAE (Jin et al., 2018), and STONED (Nigam et al., 2021)). Additionally, we demonstrate GraphXForm’s unique flexibility and stability by incorporating structural constraints conceptually suited for solvent design, such as preventing certain bond types or preserving complex molecular substructures, allowing us to design solvents with highly tailored properties. This capability highlights the strength of our approach in tackling design tasks that are difficult or infeasible for existing methods.
@@
  • We introduce a graph transformer architecture that takes a molecular graph as input and outputs probability distributions for atom placement and bond modification. This approach maintains the notion of using transformers for molecular design while moving away from string-based methods.
@@
  • We demonstrate that our method outperforms state-of-the-art molecular design techniques on two solvent design tasks and can be easily adapted to satisfy structural constraints or preserve specific molecular structures.
@@ 2.1 Molecular design
We denote by $\mathcal{M^{\prime}}$ the space of all molecular graphs $m=(\bm{a},\bm{B})$ as described above. Accordingly, let $\mathcal{M}\subset\mathcal{M^{\prime}}$ be the subspace of molecular graphs that are chemically valid, i.e., all non-hydrogen atoms follow the octet rule. We refer to following the octet rule as satisfying the ’valence constraints’ throughout the paper.
@@
Similar to Zhou et al. (2019), we pose molecular design as a sequential Markov decision process (MDP), where an agent assembles a molecular graph by iteratively adding atoms or increasing the bond order between atoms. We note that in the graph, hydrogens are always only considered implicitly, and an addition of an atom or a bond leads to a replacement of an implicit hydrogen.
On a high level, a single molecule is constructed by the agent as follows: The agent observes an initial molecule $m^{(0)}$, and then performs some action $x^{(0)}$ on it, which results in a new molecule $m^{(1)}$. Then, the agent observes $m^{(1)}$, decides on an action $x^{(1)}$ leading to molecule $m^{(2)}$, and so on. This iterative design process terminates at some molecule $m^{(T)}$ once the agent decides to perform a special DontChange action, which does not alter the molecular graph but rather marks the design as finished.
We formalize this as follows: Let $m^{(t)}=(\bm{a}^{(t)},\bm{B}^{(t)})$ be some molecule with atoms $\bm{a}^{(t)}\in\{1,\dots,k\}^{n^{(t)}}$ and bond matrix $\bm{B}^{(t)}\in\mathbb{N}_{0}^{n^{(t)}\times n^{(t)}}$ as above. We transition to a new molecule $m^{(t+1)}=(\bm{a}^{(t+1)},\bm{B}^{(t+1)})$ by performing some action $x^{(t)}\in\mathcal{A}$ on $m^{(t)}$. An action $x^{(t)}$ in the action space $\mathcal{A}$ is of one of the following three types:
@@
    $x^{(t)}=\textsc{AddAtom}(j,l)$, with $j\in\{1,\dots,k\}$ and $l\in\{1,\dots,n^{(t)}\}$. This action adds an atom of type $\Sigma_{j}$ to the graph and bonds it to the $l$-th atom. In particular, we set $n^{(t+1)}=n^{(t)}+1$. For the new molecule $m^{(t+1)}=(\bm{a}^{(t+1)},\bm{B}^{(t+1)})$, the atom vector $\bm{a}^{(t+1)}\in\mathbb{N}^{n^{(t+1)}}$ is obtained by appending $j$ to $\bm{a}^{(t)}$. The second entry $l\in\{1,\dots,n^{(t)}\}$ indicates that we bond this new $(n+1)$-th atom to the $l$-th atom, i.e., $\bm{B}^{(t+1)}\in\mathbb{N}_{0}^{n^{(t+1)}\times n^{(t+1)}}$ is obtained by appending a zero row and column to $\bm{B}^{(t)}$ and setting $B^{(t+1)}_{n+1,l}=B^{(t+1)}_{l,n+1}=1$.
@@
    $x^{(t)}=\textsc{IncBond}(j,l)$, meaning that we are increasing the bond order between atoms $j$ and $l$. In particular, we have $n^{(t+1)}=n^{(t)}$ and $\bm{a}^{(t+1)}=\bm{a}^{(t)}$. The bond matrix $\bm{B}^{(t+1)}$ is obtained from $\bm{B}^{(t)}$ by setting $B^{(t+1)}_{j,l}=B^{(t+1)}_{l,j}=B^{(t)}_{j,l}+1$. Note that this allows not only higher order bonds, but also bonds between previously unbonded atoms.
Given a molecule $m^{(0)}$ and a sequence of actions $x^{(0)},\dots,x^{(t-1)}$, we will also write $(m^{(0)},x^{(0)},\dots,x^{(t-1)})$ for the molecule $m^{(t)}\in\mathcal{M}$ that results from $m^{(0)}$ by sequentially applying the actions $x^{(0)},\dots,x^{(t-1)}$ to $m^{(0)}$. We note that, in general, the action sequence $x^{(0)},\dots,x^{(t-1)}$ is not unique to from $m^{(0)}$ to $m^{(t)}$.
Starting from a molecule in $\mathcal{M}$, by removing chemically invalid actions from the action space (which would lead to violation of valence constraints) at each step, we can guarantee to stay in the space of chemically valid molecules. Once the agent chooses the DontChange action, the design process is considered finished. We note that, starting from an appropriate initial atom, it is in fact possible to reach every target molecule in the chemically valid space $\mathcal{M}$ (i.e., all molecular graphs that can be constructed from the alphabet $\Sigma$ and that satisfy the valence constraints) when starting from any atom that exists in the target molecule. For an illustrative molecule construction, see Fig. 1.
Each addition of a bond or atom removes an implicit hydrogen from the respective atom, allowing for a straightforward representation of ionization states: Ions can be incorporated simply by adding charged atom types to the predefined alphabet, without the need for additional modifications to the action space. In contrast, the graph-based approach can not handle stereochemistry in a straightforward way.
Figure 1: Example for the sequential application of actions $x^{(0)},\dots,x^{(3)}$ to a molecule $m^{(0)}$, using the alphabet $\Sigma=(\mathrm{C},\mathrm{N},\mathrm{O})$. We show the index of each atom, which can be arbitrarily chosen at the beginning. Light blue indicates where in the graph an action is applied. The last action is DontChange, which does not change the molecular graph, but marks it as a finished design.
@@
Our graph-based approach allows us to extend this concept of action masking to enforce additional constraints, such as particular bonding patterns, minimum/maximum number of atoms and their types, or the presence of structural motifs like rings. We explore these constraints in detail in Section 3. We note that it is possible to simply assign an objective value of $\infty$- ∞ to molecules that violate these constraints after the design process. However, the ability to preemptively avoid constraint-violating regions by masking actions during the design process can significantly reduce the search space. In contrast, methods that generate molecules by sequentially synthesizing SMILES (Gupta et al., 2018; Xu et al., 2024; Mazuz et al., 2023) strings struggle to enforce such constraints during the construction process while also adhering to the grammatical rules of SMILES.
@@ 2.2 Learning algorithm
We now introduce our method for training the policy neural network $\pi_{\theta}$. Our proposed learning algorithm is a hybrid of the deep cross-entropy method (CEM) (Wagner, 2021) and the self-improvement learning (SIL) (Huang et al., 2023; Corsini et al., 2024; Pirnay & Grimm, 2024a, b) strategy. Both approaches train the network over multiple epochs in a self-improving loop, where, in each epoch, the current policy is used to generate a set of action sequences. The best sequences are selected as ’pseudo-labels’, i.e., they serve as the dataset for supervised training. The network is then trained for one epoch to assign higher probabilities to these sequences before the process is repeated with the updated network to generate new action sequences. Unlike other deep RL algorithms, this approach does not require reward shaping or value approximation. Furthermore, training in a supervised way provides stability and allows for the use of larger, decoder-only transformer architectures (Vaswani, 2017), which would be prohibitively slow under standard RL. However, SIL and the deep CEM differ in key aspects: The deep CEM is formulated for problems with a single instance (as in our molecular design task) and retains a certain percentage of the best action sequences in each epoch, which are obtained through simple sampling from the policy’s predicted distributions at each step. SIL, on the other hand, is typically applied to problems with infinitely many instances and employs more advanced sequence decoding techniques, such as sequence sampling without replacement (Kool et al., 2019), to improve solution quality and diversity while maintaining efficient decoding speed.
In our proposed approach, we aim to develop a method that works for single-instance problems, as in the deep CEM, while retaining the diverse sampling mechanism used in SIL methods (Pirnay & Grimm, 2024a, b). The pseudocode for our approach is presented in Algorithm 1, and we describe the key steps below:
@@
    Throughout the training process, we maintain a set BestFound of the best molecules discovered, which will be returned at the end. Initially, this set contains only the starting molecule $m_{0}$.
@@
    In each epoch, we incrementally sample action sequences without replacement from the policy until either an improved molecule is found or a maximum number of samples is reached (lines 5-10). For sampling sequences without replacement, we use the efficient batch-wise incremental method introduced by Shi et al. (2020), which relates to Stochastic Beam Search (Kool et al., 2019) and is employed in SIL methods (Pirnay & Grimm, 2024a, b). Sampling without replacement ensures that unique action sequences are generated, exploring the search space effectively (especially for shorter action sequences (Shi et al., 2020)). We emphasize that, while different action sequences may lead to the same molecule, this does not pose an issue because the policy is only concerned with generating the action sequences rather than the molecular structure. Since the goal of sampling sequences without replacement is to produce distinct sequences, the policy still generates diverse outputs, even if the resulting molecules can be identical.
@@
    After sampling, the set BestFound serves as a training dataset (lines 12-14). Similar to how decoder-only models in language modeling are trained to predict the next token from partial text, we sample batches of intermediate molecules and train the network with a cross-entropy loss to predict the next action for the corresponding molecule.
@@
Input: $q_{\min},q_{\max}\in\mathbb{N}$: minimum and maximum number of molecules to sample without replacement

### v2

@@ Abstract
Generative deep learning has become pivotal in molecular design for drug discovery, materials science, and chemical engineering. A widely used paradigm is to pretrain neural networks on string representations of molecules and fine-tune them using reinforcement learning on specific objectives. However, string-based models face challenges in ensuring chemical validity and enforcing structural constraints like the presence of specific substructures. We propose to instead combine graph-based molecular representations, which can naturally ensure chemical validity, with transformer architectures, which are highly expressive and capable of modeling long-range dependencies between atoms. Our approach iteratively modifies a molecular graph by adding atoms and bonds, which ensures chemical validity and facilitates the incorporation of structural constraints. We present GraphXForm, a decoder-only graph transformer architecture, which is pretrained on existing compounds and then fine-tuned using a new training algorithm that combines elements of the deep cross-entropy method and self-improvement learning. We evaluate GraphXForm on various drug design tasks, demonstrating superior objective scores compared to state-of-the-art molecular design approaches. Furthermore, we apply GraphXForm to two solvent design tasks for liquid–liquid extraction, again outperforming alternative methods while flexibly enforcing structural constraints or initiating design from existing molecular structures.
Machine Learning, Molecule Design
@@ 1 Introduction
To address the challenge of capturing long-range dependencies in sequential data that RNNs face, transformers (Vaswani, 2017) have been successfully applied due to their ability to model long-range dependencies more efficiently and as they are widely used in language modeling today (Xu et al., 2024). However, transformers are resource-intensive, and fine-tuning them with RL further constrains the model sizes that are practical in real-world applications (Xu et al., 2024; Henderson et al., 2018; Schulman et al., 2017). In general, chemical language models may propose string representations of molecules with invalid chemical structures – for example, when SMILES syntax or valence constraints are violated – which has led to numerous works aimed at circumventing this problem (O’Boyle & Dalke, 2018; Krenn et al., 2020; Cheng et al., 2023; Dai et al., 2018). Despite recent evidence that invalid SMILES can actually be beneficial from a language modeling perspective (Skinnider, 2024), they can harm the RL component of the pipeline as they can increase sample complexity and necessitate reward shaping to account for them. Moreover, the sequential nature of string synthesis makes it challenging to enforce structural constraints – such as ensuring a minimum number of specific atom types, restricting bonding patterns, or initiating design from predefined molecular substructures – often requiring scaffold-constrained techniques (Langevin et al., 2020; Arús-Pous et al., 2020).
@@
An alternative approach is to represent molecules directly as graphs, where atoms are nodes and bonds are edges, and to develop models that modify a molecular graph or design it directly (Wang et al., 2023; Li et al., 2018; Jin et al., 2018; Zhou et al., 2019; Jensen, 2019; De Cao & Kipf, 2018; Zang & Wang, 2020; Zhang et al., 2023b; Mahmood et al., 2021; Maziarka et al., 2019; Rittig et al., 2023d). Working at the graph level allows explicit encoding of atomic interactions and bonding rules, ensuring chemical validity, and makes it straightforward to start from existing structures and modify them. For example, graph-based methods like GraphGA (Jensen, 2019) employ genetic algorithms to modify molecular graphs directly, even outperforming several neural-based string-synthesis methods without relying on neural networks. As a deep learning example, Zhou et al. (2019) use deep Q-Learning (Mnih, 2013) with a simple feedforward network to optimize graphs from scratch by adding or removing atoms and bonds.
@@
We propose to combine and extend the strengths of both paradigms: employing transformers for their ability to capture long-range dependencies in sequence data, and leveraging pretraining on existing molecules, all while working directly on the molecular graph. More specifically, we use graph transformers (Ying et al., 2021; Maziarka et al., 2019) and formulate molecule design as a sequential task, where a molecular graph – starting from an arbitrary structure – is iteratively extended by placing atoms and adding bonds.
@@
To this end, we introduce GraphXForm, a decoder-only graph transformer architecture that guides these incremental decisions, predicting the next modification based on the current molecular graph. Pretrained on existing compounds, we propose a fine-tuning approach for downstream tasks that combines elements of the deep cross-entropy method (Wagner, 2021) and self-improvement learning (Pirnay & Grimm, 2024b). Unlike commonly used deep RL methods like REINFORCE (Williams, 1992), this approach allows for stable training of deep transformers with many layers.
@@
We test GraphXForm for two molecular applications: (1) drug development and (2) solvent design. While algorithmic advances in molecular design have been primarily focused on de novo drug development and corresponding benchmarks (Brown et al., 2019; Polykovskiy et al., 2020; Gao et al., 2022; Thomas et al., 2024), other areas such as materials science and chemical engineering have recently gained attention. Recent applications include catalyst (Schilter et al., 2023), fuel (Rittig et al., 2023d; Sarathy & Eraqi, 2024), polymer (Jiang et al., 2024; Yue et al., 2025), surfactant (Nnadili et al., 2024), chemical reaction substrate (Nigam et al., 2023), and solvent (König-Mattern et al., 2024) design. Thus, we test GraphXForm in both established and newer application areas.
First, we consider drug design by evaluating GraphXForm on the goal-directed tasks of the well-established GuacaMol benchmark (Brown et al., 2019). The benchmark includes multiple design tasks such as drug rediscovery, isomer identification, and multi-property optimization. Based on these different design objectives, we demonstrate the soundness of GraphXForm and its competitive performance with state-of-the-art molecular design approaches.
@@
Secondly, we apply GraphXForm for the design of solvents, which play a vital role in industrial chemical processes such as reactions, separations, and extractions. While König-Mattern (König-Mattern et al., 2024) recently applied a graph-based genetic algorithm for solvent design, the use of generative ML-based approaches remains underexplored. Therefore, our goal is to compare newer design methods and expand their capabilities by considering molecular structure constraints. Specifically, we evaluate GraphXForm on two liquid-liquid extraction tasks. The objective function in these tasks is defined by a separation factor based on activity coefficients at infinite dilution. For both downstream tasks, we compare GraphXForm to state-of-the-art molecular design approaches (Graph GA (Jensen, 2019), REINVENT-Transformer (Xu et al., 2024), Junction Tree VAE (Jin et al., 2018), and STONED (Nigam et al., 2021)). Additionally, we demonstrate GraphXForm’s flexibility and stability by incorporating structural constraints conceptually suited for solvent design, such as preventing certain bonding patterns or preserving molecular substructures, allowing GraphXForm to propose design candidates with highly tailored properties. This capability highlights the strength of our approach in tackling design tasks that are difficult for existing methods.
@@
  • We introduce a graph transformer architecture that takes a molecular graph as input and outputs probability distributions for atom and bond placement. This approach maintains the notion of using transformers for molecular design while moving away from string-based methods.
@@
  • We demonstrate that our method outperforms state-of-the-art molecular design techniques on well-established drug design benchmarks and two solvent design tasks.
  • We show that our method can be easily adapted to meet structural constraints by preserving or excluding specific molecular moieties and starting the design from initial structures.
@@ 2.1 Molecular design
We denote by $\mathcal{M^{\prime}}$ the space of all molecular graphs $m=(\bm{a},\bm{B})$ as described above. Accordingly, let $\mathcal{M}\subset\mathcal{M^{\prime}}$ be the subspace of molecular graphs that are chemically valid, that is, all non-hydrogen atoms follow the octet rule. We refer to following the octet rule as satisfying the ’valence constraints’ throughout the paper. In this framework, ionization can be incorporated straightforwardly by extending the alphabet with additional atom types that allow for an adjusted number of non-hydrogen bonds. For instance, to model ionized carbon atoms, we can include the types $\mathrm{C^{+}}$ and $\mathrm{C^{-}}$ in $\Sigma$, which are designated with a positive or negative charge and are permitted to form up to five and three non-hydrogen bonds, respectively. Chirality can be handled in an analogous manner. We refer to the Supplementary Information for the complete alphabet used.
@@
Similar to Zhou et al. (2019), we pose the molecular design as a sequential Markov decision process (MDP), where an agent assembles a molecular graph by iteratively adding atoms or bonds between atoms. We note that in the graph, hydrogens are always only considered implicitly, and an addition of an atom or a bond leads to a replacement of implicit hydrogen.
On a high level, a single molecule is constructed by the agent as follows: The agent observes an initial molecule $m^{(0)}$ and then performs some action $x^{(0)}$ on it, resulting in a new molecule $m^{(1)}$. Then, the agent observes $m^{(1)}$, decides on an action $x^{(1)}$ leading to molecule $m^{(2)}$, and so on. This iterative design process ends on some molecule $m^{(T)}$ once the agent decides to perform a special action DontChange, which does not alter the molecular graph but rather marks the design as completed.
We formalize this as follows. Let $m^{(t)}=(\bm{a}^{(t)},\bm{B}^{(t)})$ be some molecule with atoms $\bm{a}^{(t)}\in\{1,\dots,k\}^{n^{(t)}}$ and bond matrix $\bm{B}^{(t)}\in\mathbb{N}_{0}^{n^{(t)}\times n^{(t)}}$ as above. We transition to a new molecule $m^{(t+1)}=(\bm{a}^{(t+1)},\bm{B}^{(t+1)})$ by performing some action $x^{(t)}\in\mathcal{A}$ on $m^{(t)}$. An action $x^{(t)}$ in the action space $\mathcal{A}$ is of one of the following three types:
@@
    $x^{(t)}=\textsc{AddAtom}(j,l,o)$, with $j\in\{1,\dots,k\}$, $l\in\{1,\dots,n^{(t)}\}$, and $o\in\mathbb{N}$. This action adds an atom of type $\Sigma_{j}$ to the graph and connects it to the $l$-th atom with a bond of order $o$. In particular, we set $n^{(t+1)}=n^{(t)}+1$. For the new molecule $m^{(t+1)}=(\bm{a}^{(t+1)},\bm{B}^{(t+1)})$, the atom vector $\bm{a}^{(t+1)}\in\mathbb{N}^{n^{(t+1)}}$ is obtained by appending $j$ to $\bm{a}^{(t)}$. The second entry $l\in\{1,\dots,n^{(t)}\}$ indicates that we bond this new $(n+1)$-th atom to the $l$-th atom with order $o$, i.e., $\bm{B}^{(t+1)}\in\mathbb{N}_{0}^{n^{(t+1)}\times n^{(t+1)}}$ is obtained by appending a zero row and column to $\bm{B}^{(t)}$ and setting $B^{(t+1)}_{n+1,l}=B^{(t+1)}_{l,n+1}=o$.
@@
    $x^{(t)}=\textsc{AddBond}(j,l,o)$, meaning that we are adding a bond of order $o\in\mathbb{N}$ between existing, unbonded atoms $j$ and $l$. In particular, we have $n^{(t+1)}=n^{(t)}$ and $\bm{a}^{(t+1)}=\bm{a}^{(t)}$. The bond matrix $\bm{B}^{(t+1)}$ is obtained from $\bm{B}^{(t)}$ by setting $B^{(t+1)}_{j,l}=B^{(t+1)}_{l,j}=o$.
Given a molecule $m^{(0)}$ and a sequence of actions $x^{(0)},\dots,x^{(t-1)}$, we will also write $(m^{(0)},x^{(0)},\dots,x^{(t-1)})$ for the molecule $m^{(t)}\in\mathcal{M}$ that results from $m^{(0)}$ by sequentially applying the actions $x^{(0)},\dots,x^{(t-1)}$ to $m^{(0)}$. In general, the action sequence $x^{(0)},\dots,x^{(t-1)}$ is not unique to get from $m^{(0)}$ to $m^{(t)}$.
Starting from a molecule in $\mathcal{M}$, by removing chemically invalid actions from the action space (which would lead to violation of valence constraints) at each step, we can guarantee staying in the space of chemically valid molecules. Once the agent chooses the action DontChange, the design process is considered complete. We note that starting from an appropriate initial atom, it is in fact possible to reach every target molecule in the chemically valid space $\mathcal{M}$ (i.e., all molecular graphs that can be constructed from the alphabet $\Sigma$ and that satisfy the valence constraints) when starting from any atom that exists in the target molecule. For an illustrative molecule construction, see Fig. 1.
Figure 1: Example for the sequential application of actions $x^{(0)},x^{(1)},x^{(2)}$ to a molecule $m^{(0)}$, using the alphabet $\Sigma=(\mathrm{C},\mathrm{N},\mathrm{O})$. We show the index of each atom, which can be arbitrarily chosen at the beginning. Light blue indicates where in the graph an action is applied. The last action is DontChange, which does not change the molecular graph, but marks it as a complete design.
@@
Our graph-based approach allows us to extend this concept of action masking to enforce additional constraints, such as particular bonding patterns, minimum/maximum number of atoms and their types, or restricting structural motifs like rings. We explore these constraints in detail in Section 3. We note that it is possible to simply assign an objective value of $\infty$- ∞ to molecules that violate these constraints after the design process. However, the ability to preemptively avoid constraint-violating regions by masking actions during the design process reduces the search space.
@@ 2.2 Learning algorithm
We now introduce our method for training the policy neural network $\pi_{\theta}$. Our proposed learning algorithm is a hybrid of the deep cross-entropy method (CEM) (Wagner, 2021) and self-improvement learning (SIL) (Huang et al., 2023; Corsini et al., 2024; Pirnay & Grimm, 2024a, b), a sampling-based approach to expert iteration (Anthony et al., 2017). Both the deep CEM and SIL train the neural network over multiple epochs in a self-improving loop. In each epoch, the current policy is used to generate a set of action sequences, from which the best sequences are selected as ’pseudo-labels’ to serve as the dataset for supervised training. The network is then trained for one epoch to assign higher probabilities to these sequences, and the updated network is subsequently used to generate new action sequences. Unlike many deep RL algorithms, this approach does not require reward shaping or value approximation. Moreover, training in a supervised way provides stability and facilitates the use of larger, decoder-only transformer architectures (Vaswani, 2017).
There are key differences between SIL and the deep CEM. The deep CEM is formulated for problems with a single instance (as in our molecular design task) and retains a fixed percentage of the best action sequences in each epoch, which are obtained through simple sampling from the policy’s predicted distributions at each step. In contrast, SIL is typically applied to problems with infinitely many instances and employs more advanced sequence decoding techniques – such as sequence sampling without replacement (Kool et al., 2019) – to improve solution quality and diversity while maintaining efficient decoding speed.
In our proposed approach, we seek to develop a method that works for single-instance problems, as in the deep CEM, while retaining the diverse sampling mechanism used in SIL methods (Pirnay & Grimm, 2024a, b). In particular, we adopt the Take a Step and Reconsider (TASAR) method (Pirnay & Grimm, 2024b) to sample action sequences, as detailed below. The pseudocode for our approach is presented in Algorithm 1, and we describe the key steps as follows:
@@
    Throughout training, we maintain a set BestFound containing the best molecules discovered so far. This set is initially initialized with the starting molecule $m_{0}$.
@@
    In each epoch, we sample action sequences from the policy using the TASAR method (Pirnay & Grimm, 2024b). Specifically, $\beta$ action sequences are sampled without replacement using Stochastic Beam Search (Kool et al., 2019) with a beam width $\beta\in\mathbb{N}$. These sequences are then evaluated using the objective function. The best action sequence among the $\beta$ samples is followed for a predefined number $\sigma$ of actions. Subsequently, alternative, previously unseen actions are sampled from the resulting partial sequence to further explore and potentially improve the solution, and the process continues. Sampling without replacement ensures that unique action sequences are generated, effectively exploring the search space (particularly for shorter sequences (Shi et al., 2020)). Although different action sequences may result in the same molecule, this does not pose an issue for TASAR since the policy is concerned with generating action sequences rather than the molecular structure itself. By sampling sequences without replacement, the policy is encouraged to produce diverse outputs, even if the resulting molecules are identical.
@@
    After sampling, the set BestFound serves as the training dataset for supervised learning (lines 6-8). Similarly to how decoder-only models in language modeling are trained to predict the next token from partial text, we sample batches of intermediate molecules and train the network with a cross-entropy loss to predict the next action for the corresponding molecule.
@@
Input: $\beta,\sigma\in\mathbb{N}$: beam width and step size (hyperparams for TASAR (Pirnay & Grimm, 2024b))